OpenStudy (anonymous):

HELP PLEASE |3x-9| = 0 what are 2 equations in this?!

OpenStudy (texaschic101):

| 3x - 9 | = 0 3x - 9 = 0 and -(3x - 9) = 0

OpenStudy (anonymous):

Thanks my teacher asked me im like i have no clue. Thanks alot though

OpenStudy (texaschic101):

no problem :)